OpenStudy (anonymous):

Need help with finding equation of given circle center (3, 3) radius of 2 How do I solve this?

OpenStudy (lucaz):

take a look here http://www.purplemath.com/modules/sqrcircle.htm

OpenStudy (lucaz):

you already has the center and the radius, you have to do steps backwards

OpenStudy (anonymous):

can you walk me through this

OpenStudy (lucaz):

one sec

OpenStudy (anonymous):

i dont know which format to even use

OpenStudy (whpalmer4):

formula for a circle with center at (h,k) and radius r is (x-h)^2 + (y-k)^2 = r^2 plug in the numbers...

OpenStudy (anonymous):

thx

OpenStudy (whpalmer4):

you have center of (3,3) = (h,k) so h = 3, k = 3 radius = 2 so r = 2

OpenStudy (anonymous):

so you can use both of those formula's?

OpenStudy (lucaz):

you can expand the quation to get ax^2+by^2+cx+dy=e

OpenStudy (anonymous):

is that what you did?

OpenStudy (lucaz):

it depends on what information you're given

OpenStudy (whpalmer4):

normally, I'd leave it in the (x-h)^2 + (y-k)^2 = r^2 form unless specifically instructed to expand it. You can easily see the details of the circle in the former form, which is useful. The latter form you can't see diddly unless you do the algebra to convert it back.

OpenStudy (lucaz):

yeah, I agree

OpenStudy (anonymous):

ok so (x-h)^2+(y-k)^2 after I plug in those numbers what do i do? I mean it'll be just like x-3+y+3=2

OpenStudy (anonymous):

and I can't do anything with x-3? can I except combine them

OpenStudy (whpalmer4):

no, it will be (x-3)^2 + (y-3)^2 = 2^2

OpenStudy (anonymous):

Yeah I was just using for example

OpenStudy (anonymous):

ok nvm

OpenStudy (whpalmer4):

but you didn't have any exponents in there

OpenStudy (whpalmer4):

you have to have the x term squared, and the y term squared, or it won't be a circle.

OpenStudy (anonymous):

Im using (x-h)^2+(y-k)^2 right?

OpenStudy (whpalmer4):

yes. notice that if the center of the circle is at (0,0) that equation reduces to (x-0)^2 + (y-0)^2 = r^2 x^2 + y^2 = r^2 which is just the Pythagorean theorem in disguise!

OpenStudy (anonymous):

ok hold on let me do this and see if I understand give me a min

OpenStudy (anonymous):

What do I plug in for k? 3?

OpenStudy (whpalmer4):

if the draw tool worked, I'd draw you a picture but the equation for the circle is just showing you all the points that are the same distance away from the center. the distance formula is also the Pythagorean theorem in disguise: d = sqrt((x2-x1)^2 + (y2-y1)^2) d^2 = (x2-x1)^2 + (y2-y1)^2 (formula for distance between (x1, y1) and (x2, y2))

OpenStudy (whpalmer4):

our center is at (3,3), right? so the formula expects the center at (h,k) so h = 3, k = 3

OpenStudy (anonymous):

ah

OpenStudy (anonymous):

Yes if the draw tool was working I'd understand much better

OpenStudy (anonymous):

Im more of a visual learner

OpenStudy (anonymous):

let me plug it in

OpenStudy (whpalmer4):

it unnecessarily confuses things to use the same number, in my opinion. I would have written a problem where the numbers were all different, just to make it that much clearer which bit goes where.

OpenStudy (anonymous):

x-3^2+y-3^2=2^2 correct?

OpenStudy (anonymous):

>>x-6+y-6=4

OpenStudy (anonymous):

after simplifying it

OpenStudy (anonymous):

well centre is 3,3 and radius is 2. equations of a circle is x^2 + y^2 = r^2

OpenStudy (whpalmer4):

no! you MUST have the exponents and parentheses! (x-3)^2 + (y-3)^2 = 2^2

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

Yeah I know

OpenStudy (anonymous):

(x-6)+(y-6)=4

OpenStudy (whpalmer4):

(x-3)^2 + (y-3)^2 = 2^2 (x^2 -3x - 3x + 9) + (y^2 -3y -3y + 9) = 4 x^2 - 6x + 9 + y^2 - 6y +9 = 4 x^2 - 6x + y^2 - 6y + 18 = 4 x^2 - 6x + y^2 - 6y = -14 sorry @dg98 the "NO" was directed at Mr. Ick

OpenStudy (anonymous):

is there no way You can draw this out?

OpenStudy (whpalmer4):

@Mr.Ick stop skipping the ^2 !!!

OpenStudy (anonymous):

hold on

OpenStudy (whpalmer4):

Sorry, I have to leave, but it's all there. Just read through my posts again.

OpenStudy (anonymous):

thats not right

OpenStudy (whpalmer4):

what's not right? the equation of a circle with center at (3,3) and radius 2 is (x-3)^2 + (y-3)^2 = 2^2 or (x-3)^2 + (y-3)^2 = 4 if you expand it, you get x^2 + y^2 - 6x - 6y + 18 = 4 or x^2 + y^2 - 6x - 6y + 14 = 0 here's a picture: http://www.wolframalpha.com/input/?t=crmtb01&f=ob&i=circle%20with%20center%20(3%2C3)%20and%20radius%202